• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 16:02:31 on September14,2005

    Entry number 154727

    keyword=day summary

    Day shift with LeRose and Michaels
    
    Production running.
    
    IHWP .... run .... good pairs
    
    OUT ..... 5041 .... 49 K
    OUT ..... 5042 .... 43 K
    OUT ..... 5043 .... 26 K
    
    Helicity Magnet Tests -- S. Baily made halog entry 154713 about
    the detailed conditions.  Runs 5044 through 5048.
    
    10:40 Helicity magnet turned off.  Return to normal production.
    Run flipper, restart Compton
    
    IN ..... 5049 ..... 42 K
    IN ..... 5050 ..... 45 K
    IN ..... 5051 ..... ended because of ACC problems.
    
    No beam the rest of the shift.
    We wanted to do a Q2 scan, but didn't have the chance.